OpenStudy (anonymous):

∫[(4x^3)dx/√3:(x^4-9)]=

OpenStudy (anonymous):

put x^4-9= t, (4x^3)dx=dt now equation become int dt/3t. (logt)/3={ log(x^4-9)}/3